# Repeatable Energy-Efficient Perching for Flapping-Wing Robots Using Soft Grippers

### v1

@@ 2.3 Assembly and Integration
In addition to gaining complete landing, perching and take-off functionality and seeing optimisations regarding the use in FWMAVs, the gripping mechanism has been made lighter compared to the previous proof of concept, weighing 38.8 g instead of 45 g. Furthermore, the mechanism features a more compact design, occupying a smaller footprint and containing thinner grippers while exerting larger gripping forces.
This ensures usage in small- to medium-sized FWMAVs that are capable of handling at least this amount of payload. The gripping power mainly constricts the upper size limit of an applicable FWMAV and can, therefore, be defined in different ways. Furthermore, this limit can be easily shifted upward using slightly heavier but increasingly more powerful servo motors.
As shown in the weight distribution diagram (see Fig. 4), the active-passive actuation system takes up the largest part at 44 %, dominated by the heavy but powerful servo-motor. The ratchet mechanism enables energy-efficient and prolonged perching but only inflicts a negligible weight penalty. The remaining 56 % are almost evenly divided by the mechanical linkage (29 %) and the flexible soft grippers (27 %). Would this system be upscaled to fit larger sizes of flapping-wing MAVs, requiring higher gripping strengths, the actuation system would scale favourably and decrease in prominence due to more capable servo-motors.
In order to interface with the active-passive perching system and handle gripping and release states, a power supply and control board had been designed for the first prototype (Broers and Armanini,, 2022). Actuation was controlled by a microcontroller that steers the servo motor and can cut off its power through a MOSFET circuit. For testing purposes and to guarantee modularity for different testing platforms, the circuitry was established as a simple breadboard and not integrated into the platform’s onboard electronics.
@@
To demonstrate the functionality of the perching system and test its capabilities and suitability for FWMAV, a representative and commercially available drone was chosen with the “Flapper Nimble+” (see Fig. 1), made by Flapper Drones, (2024). The Flapper is a lightweight and tailless flapping wing MAV with hover capabilities, enabling a wide variety of flight manoeuvres. Its four wings, consisting of two pairs, are arranged in an X-configuration with a wingspan of 490 mm. The flapping motion for each wing pair is created by a brushless DC-motor-powered gear train and linkage, resulting in a flapping frequency of 20 Hz at full throttle (Flapper Drones,, 2024). Manoeuvres are achieved through wing geometry and wing motion alterations, which are enabled by two servo motors and two independent brushless DC motors. As the tailless configuration makes the Nimble+ inherently unstable, a Crazyflie Bolt 1.1 flight controller is used. The complete system is powered by a single 300 mAh 2S LiPo battery, enabling a flight time of more than five minutes hovering with a payload of 25 g (Flapper Drones,, 2024).
For integration into the proposed system, the Flapper Nimble+ was used in its bare configuration (i.e. without the attachable landing gear, body shells, and head shell) to minimise take-off weight. In addition, the control board was affixed to its carbon fibre spine, which was opposite the battery. The gripping mechanism was then connected to the bottom of the spine by inserting the squared spine into the counterpart fixture on the top bar of the linkage. The power supply of the Flapper was connected to the control board and Flapper drone via a split connector. Lastly, the gripping mechanism was attached to the board with small detachable connectors, allowing for quick exchange during maintenance or a change of component position.
After integration, this puts the combined system’s weight, including the flapper drone (89 g), battery (18 g), gripping mechanism (38.8 g), and control board (26 g), at roughly 172 g of take-off mass.
@@ 3 Testing and Results
Three different stages of tests were conducted after the functional prototype was completed and integrated onto the Flapper drone. Given that the performance of different gripper configurations was already tested in our previous work, this set of tests concentrated on the performance of the full, integrated system.
The first stage consisted of static tests, in which the system’s gripping performance was assessed against disturbance forces and moments of different magnitudes and orientations.
To provide a complete and comprehensive evaluation of the system, we tested all three disturbance force directions, as well as the resulting disturbing moments that act on the integrated system.
Additionally, we also accounted for a special “upside-down” perched case and different perching-object inclinations to get a fuller understanding of the robustness of the system while perched. Considering different inclinations provides an overview of the set of perching conditions that can successfully be handled by our mechanism, and is essential because in a real-world setting branches or pipes will often not be perfectly aligned with either the ground or the robot. The “upside-down” perching case is important because our perching mechanism, like many others, is attached to the lower end of the robot, while the CG of our robot is relatively high up: therefore the perched robot is particularly susceptible to transversal forces, which may make it lose its balance and topple over. However, if the gripping force is such that the robot remains in a hanging, bat-like position, rather than drop off the perch, the rolling over motion becomes less problematic – given sufficient space, the robot is expected to be able to take off from an upside-down position.
@@
Lastly, free-flight tests under realistic conditions were carried out, demonstrating the complete perching cycle of landing, holding on, and take-off.
@@ 3.1 Static Performance Testing
The testing setup, which can be seen in Fig. 5, consisted of a fixture in which different perching objects were placed and oriented at a desired inclination angle (see also Fig. 9). To repeatably and reliably assess perching performance, a variety of perches were used for these tests, i.e.: two circular PVC pipes, with diameters of 50 mm and 40 mm respectively, two circular wooden bars, with diameters of 40 mm and 30 mm, respectively, and a square wooden bar, with a cross-section of 30 mm by 30 mm, which was used both in a flat square orientation and a diamond orientation (see Fig. 5).
These objects were selected to represent common indoor and outdoor landing structures, like tree branches, railings, and struts.
A force gauge was connected to fixtures at specific positions of the drone to measure performance. Four different measuring directions were used to represent disturbance forces or torques for each degree of freedom (see coloured arrows in Fig. 5). The first measurement was taken horizontally in a perpendicular manner to the perch at a fixed distance from the pivot point, in order to measure the rotational momentum around the perch (yellow arrow). The second measurement was taken axially to the perch, again at a fixed distance to the pivot point, in order to measure the axial moment of possible disturbances (red arrow). In this way, the peak force was measured until the system’s slippage point. The remaining two measurements were taken vertically, perpendicularly to the perch, once measuring upwards with the system sitting atop the perching object (blue arrow) and once downwards, with the system “hanging” below the perch. The results of these perpendicular measurements represent the maximal force required to loosen the grip from the perch, a good indicator of general gripper strength.
@@
The results of these tests are shown in Fig. 6, which presents the resulting forces and moments in each of the three main measurement directions clarified in Fig. 5 (with the perpendicular down direction excluded here for clarity), for each of the considered perching objects.
The perpendicular forces needed to pull the system from its perch are consistently above 6 N, with an arithmetic mean of 7.4 N across all perching objects. The best-performing scenario was the diamond-oriented square wooden bar, which required 9.5 N of force. This can be explained by the better gripper performance when conforming to irregular shapes. Here, especially, the contact point of the object is in a good position, not too far out to the gripper tips, but acting on the middle portion, allowing the gripper to close further compared to a square-oriented object. The gripper strengths show a mean improvement of 42 % more force required in the upwards perpendicular direction, normalised with the system weight compared to the results obtained in previous work for a similar design concept (Broers and Armanini,, 2022). This gain can be attributed to the generally optimised design and especially overhauled ratcheting functionality. The rotational disturbance moment required for continuous displacement around the perch shows an increase with larger contact area and irregular shapes, with the smaller diameters of perching objects performing the worst as they move further from the primary gripper size design point. The best-performing case is again the square wooden bar in a diamond orientation, which can withstand a 12.7 Ncm moment of disturbance. The mean gripper strength, measured through its resistance to a rotational disturbance moment, increased by 63 % compared to the previous design (Broers and Armanini,, 2022). Lastly, the axial disturbance moment was fairly consistent across the different perching objects, with a mean of 10.3 Ncm of disturbance required for continuous displacement in the axial direction. Consistent with the other disturbance measurements, the diamond-oriented square wooden bar performed best, with a moment of 14.3 Ncm required for displacement.
@@
Thus in addition to the basic measurements discussed, further tests were conducted on the ø 40 mm circular wooden bar to fully explore the feasibility envelope of our mechanism. Specifically, the previous tests were repeated at three further perch inclinations: 5°, 10°, 12.5°. The highest inclination of 12.5° was selected following preliminary tests that indicated this to be close to the maximum inclination that could still successfully be handled. For these tests, the fixture of the bar was lifted vertically on one side, resulting in a defined inclination (see also Fig. 9 for orientation reference).
Results of these further tests can be seen in Fig. 7: In panel A, the perpendicular disturbance force for both the regular and the upside-down direction is plotted against the different perch inclinations. For the upwards perpendicular force, the results are fairly similar up to 10° of inclination, followed by a significant drop in performance at the highest inclination of 12.5°. For the downwards perpendicular direction, performance drops off at an inclination of 5°, where it remains stable even for higher inclinations. Of note for the upside-down case is that even for the highest tested inclination, roughly 300 g of weight are still supported. This indicates that even in a scenario where the robot loses its balance and reverses, it will not drop off the perch.
In panel B of Fig. 7, the rotational and axial disturbance moments are plotted against the perch inclinations. The rotational moment required for displacement increases somewhat for 5° of inclination, compared to the non-inclined case, but stays fairly consistent across all tested inclinations overall. However, a sharp linear decline in the measured moment is seen for the axial disturbance moment, nearing zero at the highest inclination setting. This is in line with expectations, as the highest inclination attainable is defined by the system’s robustness against axial moment, which increases with perch inclination as the distance between the pivot point of the system and its centre of gravity increases.
Based on these results, attachment strength measured through the perpendicular disturbance force is well-performing at inclinations up to 10°. This would enable the system to find a large variety of perching opportunities on natural structures such as inclined tree branches and man-made struts. However, the resistance against axial disturbance forces after successful perching attempts is severely diminished at these higher inclinations. Therefore, if robustness against axial disturbances is required, e.g. in windy outdoor scenarios, perching objects of lower inclination will be strongly preferred. The full feasibility envelope is further explored in testing in the following section.
@@ 3.2 Free-Fall Testing
The flight controller’s IMU provided acceleration data at 250 Hz to an onboard microSD card. Again, the Z-acceleration, i.e. vertical acceleration, was of primary interest. With the identification of the two relevant zero-crossings before the peak of maximum acceleration (clearly seen in Fig. 8.C), the velocity at impact was calculated via integration.
@@
The first test run consisted of over 26 recordings at different approach speeds, utilising both data sources simultaneously. These revealed that the Vicon and IMU data yielded similar and plausible results for the calculated impact velocities, with a mean velocity discrepancy of 6 % and a standard deviation of 5 %. In Extension 1.1, a table of these data recordings can be accessed. In Fig. 8, three distinct approaches are shown from the free-fall testing. The figure contains Vicon-measured height data (panel A), Z-velocity calculated from the Vicon position data (panel B), and IMU-measured Z-acceleration data (panel C). The Vicon and IMU datasets are synchronised from the time of impact, which was set at the data point corresponding to the highest approach velocity.
The three examples showcase three different impact velocities, ranging from 0.29 m/s to 1.29 m/s. While the lower and middle velocity still lead to a successful perch, the highest velocity case is an example of an unsuccessful perching attempt.
@@
Of note in panel A is the bounce-back after impact, which is mainly caused by the elastic trigger fork design, increasing in height with higher impact velocities. Likewise, in panel A the unsuccessful attempt (blue line) can be easily identified, as the re-attachment to the perch does not happen, i.e. rather than stabilise at zero (perch-level), the height continues to drop after impact. Similar effects can also be seen in panel C, where the bounce-back effect is clearly visible in the positive Z-acceleration peaks. Higher impact velocities, therefore, lead to a quicker and larger recoil of the system from the perch, at one point exceeding the gripper reaction time or, rather, leaving the designed closing area too early.
@@
As the recording of the Vicon data adds complexity to the setup and adds weight to the robot, due to the markers that need to be attached to it, all further 64 fall tests were conducted using only the on-board IMU data recording. A table containing the calculated velocities from these tests can be accessed in Extension 1.2. The aim of these tests was to determine the maximum feasible impact velocity through gradual variations of the drop height.
These tests were conducted for five different perch inclinations, i.e. 0°, 3°, 6°, 9° and 11° (see visualisation in Fig. 9), and the maximum successful impact velocity was obtained for each of these cases.
@@
At least five data points each, for successful attempts and for unsuccessful attempts, were collected per inclination setting. Results of these tests are displayed in Fig. 10 in the form of an approach sufficiency subspace, where the different categories of perching attempt outcomes (failed/successful/mixed) are shown in a plot of impact velocity over perch inclination. The green area in the diagram represents a 100 % success rate of perching attempts, while the red area represents predominantly failed perching attempts. The green and red crosses represent individual data points that fall within each category. The orange area in between represents a mixed success rate of perching attempts. Along the x-axis, the grey band shows the maximum inclination that was successfully tested in free-fall tests (just over 11°), while the dashed black line represents the maximum inclination achievable in the static tests (cf. Sec. 3.1), i.e. 12.5°. This discrepancy between both inclination limits is due to disturbances introduced during the approach and the recoil forces acting on impact.
@@
In general, the velocity at which the failure rate becomes dominant stays fairly constant, at around 1.3 m/s, across different inclinations. However, the highest feasible impact velocity that guarantees a successful attempt declines sharply at higher inclinations and has a maximum for the 6° inclination at 1.3 m/s. This maximum could be possibly explained by a small offset due to an unintended and imperfect centre of gravity location of the integrated system, where a slight inclination might lead to a better balance on the perching object, with optimum stability at around 6°. Alternatively, this could be due to a slightly larger and asymmetric gripper contact surface area due to the change in the overlapping area between grippers and the inclined perching object. For even larger inclinations, the rising axial momentum of the system becomes the driving factor and diminishes these gains again. Overall, the highest impact velocity leading to 100 % successful perching attempts, under up to 10° of inclination, is 1.1 m/s.

### v2

@@ 2.3 Assembly and Integration
In addition to gaining complete landing, perching and take-off functionality and being adapted for usage on FWMAVs, the gripping mechanism has been made lighter compared to the previous proof of concept, weighing 38.8 g instead of 45 g. Furthermore, the mechanism features a more compact design, occupying a smaller footprint and containing thinner grippers while exerting larger gripping forces.
This allows for usage on small- to medium-sized FWMAVs that are capable of handling at least this amount of payload. The gripping power mainly constrains the upper size limit of an applicable FWMAV and can, therefore, be defined in different ways. Furthermore, this limit can be easily shifted upward using slightly heavier but increasingly more powerful servo motors.
As shown in the weight distribution diagram (see figure 6), the active-passive actuation system takes up the largest part of the total weight at 44 %, dominated by the heavy but powerful servo-motor. The ratchet mechanism enables energy-efficient and prolonged perching but only inflicts a negligible weight penalty. The remaining 56 % are almost evenly divided between the mechanical linkage (29 %) and the flexible soft grippers (27 %). If this system were to be upscaled to fit larger flapping-wing MAVs, requiring higher gripping strengths, the actuation system would scale favourably and decrease in prominence due to more capable servo-motors.
In order to interface with the active-passive perching system and handle gripping and release states, a power supply and control board had been designed for the first prototype (Broers and Armanini, 2022). Actuation was controlled by a microcontroller that steered the servo motor and could cut off its power through a MOSFET circuit. For testing purposes and to guarantee modularity for different testing platforms, the circuitry was established as a simple breadboard and not integrated into the platform’s onboard electronics.
@@
To demonstrate the functionality of the perching system and test its capabilities and suitability for FWMAV, a representative commercially available drone was chosen for testing, i.e. the “Flapper Nimble+” (see figure 1) produced by Flapper Drones (2024). The Flapper is a lightweight and tailless flapping wing MAV with hover capabilities, enabling a wide variety of flight manoeuvres. Its four wings, consisting of two pairs, are arranged in an X-configuration with a wingspan of 490 mm. The flapping motion for each wing pair is created by a brushless DC-motor-powered gear train and linkage, resulting in a flapping frequency of 20 Hz at full throttle. Manoeuvres are achieved through wing geometry and wing motion alterations, which are enabled by two servo motors and two independent brushless DC motors. As the tailless configuration makes the Nimble+ inherently unstable, a Crazyflie Bolt 1.1 flight controller is used. The complete system is powered by a single 300 mAh 2S LiPo battery, enabling a flight time of more than five minutes hovering with a payload of 25 g (Flapper Drones, 2024).
For integration into the proposed system, the Flapper Nimble+ was used in its bare configuration (i.e. without the detachable landing gear, body shells, and head shell) to minimise take-off weight. In addition, the control board was affixed to its carbon fibre spine, which was opposite the battery. The gripping mechanism was then connected to the bottom of the spine by inserting the squared spine into the counterpart fixture on the top bar of the linkage. An alternative approach would be to attach the gripper system on top of the FWMAV’s spine, allowing for perching from below, with the robot hanging on to the perching object. This would result in a more stable perch and would be an interesting approach to test in future work. However, positioning the gripper at the top would shift the center of gravity upwards, introducing further instability into the already unstable flying platform.
The power supply of the Flapper was connected to the control board and the drone via a split connector. Lastly, the gripping mechanism was attached to the board with small detachable connectors, allowing for quick exchange during maintenance or a change of component position.
After integration, this puts the combined system’s weight, including the flapper drone (89 g), battery (18 g), gripping mechanism (38.8 g), and control board (26 g), at roughly 172 g of take-off mass. Note that this exceeds the recommended take-off weight, primarily due to the external control board; however, for our testing purposes, the resulting shortened flight time was acceptable.
@@ 3 Testing and Results
Three different stages of tests were conducted after the functional perching mechanism was completed and integrated onto the FWMAV. Given that the performance of different gripper configurations was already tested in our previous work, the current tests could directly concentrate on the performance of the full, integrated system.
The first stage consisted of static tests, in which the system’s gripping performance was assessed against disturbance forces and moments of different magnitudes and orientations. To provide a complete and comprehensive evaluation of the system, we tested all three disturbance force directions, as well as the resulting disturbing moments that act on the integrated system.
Additionally, we also accounted for a special “upside-down” perched case and different perching-object inclinations to get a fuller understanding of the robustness of the system while perched. Considering different inclinations provides an overview of the set of perching conditions that can successfully be handled by our mechanism and is essential because, in a real-world setting, branches, pipes or railings will often not be perfectly aligned with either the ground or the robot. The “upside-down” perching case is important because our perching mechanism, like many others, is attached to the lower end of the robot, while the CG of our robot is relatively high up: therefore the perched robot is particularly susceptible to transversal forces, which may make it lose its balance and topple over. However, if the gripping force is such that the robot remains in a hanging, bat-like position, rather than drop off the perch, the rolling over motion becomes less problematic – given sufficient space, the robot is expected to be able to take off from an upside-down position.
@@
Lastly, free-flight tests under realistic conditions were carried out, demonstrating the complete perching cycle of landing, holding on, and taking off.
@@ 3.1 Static Performance Testing
The testing setup, which can be seen in figure 7, consisted of a fixture in which different perching objects were placed and oriented at a desired inclination angle (see also figure 12). To repeatably and reliably assess perching performance, a variety of perches were used for these tests, i.e.: two circular PVC pipes, with diameters of 50 mm and 40 mm respectively, two circular wooden bars, with diameters of 40 mm and 30 mm, respectively, and a square wooden bar, with a cross-section of 30 mm by 30 mm, which was used both in a flat square orientation and a diamond orientation (see figure 7). These objects were selected to represent common indoor and outdoor landing structures, like tree branches, railings, and struts.
A force gauge was connected to fixtures at specific positions of the drone to measure performance. Four different measuring directions were used to represent disturbance forces or torques for each degree of freedom (see coloured arrows in figure 7). The first measurement was taken horizontally in a perpendicular manner to the perch at a fixed distance from the pivot point, in order to measure the rotational momentum around the perch (yellow arrow). The second measurement was taken axially to the perch, again at a fixed distance to the pivot point, in order to measure the axial moment of possible disturbances (red arrow). In this way, the peak force was measured until the system’s slippage point. The remaining two measurements were taken vertically, perpendicularly to the perch, once measuring upwards with the system sitting atop the perching object (blue arrow) and once downwards, with the system “hanging” below the perch. The results of these perpendicular measurements represent the maximal force required to loosen the grip from the perch, a good indicator of general gripper strength.
@@
The results of these tests are shown in figure 8, which presents the resulting forces and moments in each of the three main measurement directions clarified in figure 7 (with the perpendicular down direction excluded here for clarity), for each of the considered perching objects.
The perpendicular forces needed to pull the system from its perch are consistently above 6 N, with an arithmetic mean of 7.4 N across all perching objects. The best-performing scenario was the diamond-oriented square wooden bar, which required 9.5 N of force. This can be explained by the better gripper performance when conforming to irregular shapes. Here, especially, the contact point of the object is in a good position, not too far out to the gripper tips, but acting on the middle portion, allowing the gripper to close further compared to a square-oriented object. The gripper strengths show a mean improvement of 42 % more force required in the upwards perpendicular direction, normalised with the system weight compared to the results obtained in previous work for a similar design concept (Broers and Armanini, 2022). This gain can be attributed to the generally optimised design and especially the overhauled ratcheting functionality. The rotational disturbance moment required for continuous displacement around the perch shows an increase with larger contact area and irregular shapes, with the smaller diameters of perching objects performing the worst as they move further from the primary gripper size design point. The best-performing case is again the square wooden bar in a diamond orientation, which can withstand a 12.7 Ncm moment of disturbance. The mean gripper strength, measured through its resistance to a rotational disturbance moment, increased by 63 % compared to the previous design (Broers and Armanini, 2022). Lastly, the axial disturbance moment was largely consistent across the different perching objects, with a mean of 10.3 Ncm of disturbance required for continuous displacement in the axial direction. Consistent with the other disturbance measurements, the diamond-oriented square wooden bar performed best, with a moment of 14.3 Ncm required for displacement.
@@
Thus, in addition to the basic measurements discussed, further tests were conducted on the ø 40 mm circular wooden bar to fully explore the feasibility envelope of our mechanism. Specifically, the previous tests were repeated at three further inclinations of the perch axis: 5°, 10°, and 12.5°. The highest inclination of 12.5° was selected following preliminary tests that indicated this to be close to the maximum inclination that could still successfully be handled. For these tests, the fixture of the bar was lifted vertically on one side, resulting in a defined inclination of the perch axis compared to its horizontal state (see also figure 12 for orientation reference).
Results of these further tests can be seen in figure 9: In panel A, the perpendicular disturbance force for both the regular and the upside-down direction is plotted against the different perch inclinations. For the upwards perpendicular force, the results are fairly similar up to 10° of inclination, followed by a significant drop in performance at the highest inclination of 12.5°. For the downwards perpendicular direction, performance drops off at an inclination of 5°, where it remains stable even for higher inclinations. Of note for the upside-down case is that even for the highest tested inclination, roughly 300 g of weight are still supported. This indicates that even in a scenario where the robot loses its balance and reverses, it will not drop off the perch, acting as a useful fallback.
In panel B of figure 9, the rotational and axial disturbance moments are plotted against the perch inclinations. The rotational moment required for displacement increases somewhat for 5° of inclination, compared to the non-inclined case, but stays fairly consistent across all tested inclinations overall. However, a sharp linear decline in the measured moment is seen for the axial disturbance moment, nearing zero at the highest inclination setting. This is in line with expectations, as the highest inclination attainable is defined by the system’s robustness against axial moment, which increases with perch inclination as the distance between the pivot point of the system and its centre of gravity increases.
Based on these results, attachment strength measured through the perpendicular disturbance force is well-performing at inclinations up to 10°. This would enable the system to find a large variety of perching opportunities on natural structures such as inclined tree branches and man-made struts. However, the resistance against axial disturbance forces after successful perching attempts is severely diminished at these higher inclinations. Therefore, if robustness against axial disturbances is required, e.g. in windy outdoor scenarios, perching objects of lower inclination will be strongly preferred. To further increase this resistance, especially at higher inclinations, an optimisation of the gripper surface material would be required to provide better adhesion to the perching object. The full feasibility envelope is further explored in testing in the following section.
@@ 3.2 Free-Fall Testing
The flight controller’s IMU provided acceleration data at 250 Hz to an onboard microSD card. Again, the Z-acceleration, i.e. vertical acceleration, was of primary interest. With the identification of the two relevant zero-crossings before the peak of maximum acceleration (clearly seen in figure 10.C), the velocity at impact was calculated via integration.
@@
The first test run consisted of over 26 recordings at different approach speeds, utilising both data sources simultaneously. These revealed that the Vicon and IMU data yielded similar and plausible results for the calculated impact velocities, with a mean velocity discrepancy of 6 % and a standard deviation of 5 %. In Extension 1.1, a table of these data recordings can be accessed. In figure 10, three distinct approaches are shown from the free-fall testing. The figure contains Vicon-measured height data (panel A), Z-velocity calculated from the Vicon position data (panel B), and IMU-measured Z-acceleration data (panel C). The Vicon and IMU datasets are synchronised from the time of impact, which was set at the data point corresponding to the highest approach velocity. The three examples showcase three different impact velocities, ranging from 0.29 m/s to 1.29 m/s. While the lower and middle velocity still lead to a successful perch, the highest velocity case is an example of an unsuccessful perching attempt.
@@
Of note in panel A is the bounce-back after impact, which is mainly caused by the elastic trigger fork design, increasing in height with higher impact velocities. Likewise, in panel A, the unsuccessful attempt (blue line) can be easily identified, as the re-attachment to the perch does not happen, i.e. rather than stabilising at zero (perch-level), the height continues to drop after impact. Similar effects can also be seen in panel C, where the bounce-back effect is clearly visible in the positive Z-acceleration peaks. Higher impact velocities, therefore, lead to a quicker and larger recoil of the system from the perch, at one point exceeding the gripper reaction time or, rather, leaving the designed closing area too early.
In order to quantify the properties of the system after impact, we modelled it as a 1-DoF mass-spring-damper system and estimated the spring stiffness k and damping coefficient c using a linear regression with a dataset from the free fall tests (see figure 11(a)). This resulted in an estimated spring stiffness k = 235.8 N/m with a standard error of 22.2 N/m, and an estimated damping coefficient c = 2.44 Ns/m with a standard error of 0.5 Ns/m. We validated this model with another dataset of a fall test with lower impact velocity and achieved similar accuracy (see figure 11(b)).
@@
As the recording of the Vicon data adds complexity to the setup and adds weight to the robot, due to the markers that need to be attached to it, all further 64 fall tests were conducted using only the on-board IMU data recording. A table containing the calculated velocities from these tests can be accessed in Extension 1.2. The aim of these tests was to determine the maximum feasible impact velocity through gradual variations of the drop height. These tests were conducted for five different perch inclinations, i.e. 0°, 3°, 6°, 9° and 11° (see visualisation in figure 12), and the maximum successful impact velocity was obtained for each of these cases.
@@
At least five data points each, for successful attempts and for unsuccessful attempts, were collected per inclination setting. Results of these tests are displayed in figure 13 in the form of an approach sufficiency subspace, where the different categories of perching attempt outcomes (failed/successful/mixed) are shown in a plot of impact velocity over perch inclination. The green area in the diagram represents a 100 % success rate of perching attempts, while the red area represents predominantly failed perching attempts. The green and red crosses represent individual data points that fall within each category. The orange area in between represents a mixed success rate of perching attempts. Along the x-axis, the grey band shows the maximum inclination that was successfully tested in free-fall tests (just over 11°), while the dashed black line represents the maximum inclination achievable in the static tests (cf. section 3.1), i.e. 12.5°. This discrepancy between both inclination limits is due to disturbances introduced during the approach and the recoil forces acting on impact.
@@
In general, the velocity at which the failure rate becomes dominant stays fairly constant, at around 1.3 m/s, across different inclinations. However, the highest feasible impact velocity that guarantees a successful attempt declines sharply at higher inclinations and has a maximum for the 6° inclination at 1.3 m/s. This maximum could possibly be explained by a small offset due to an unintended and imperfect centre of gravity location of the integrated system, where a slight inclination might lead to a better balance on the perching object, with optimum stability at around 6°. Alternatively, this could be due to a slightly larger and asymmetric gripper contact surface area due to the change in the overlapping area between the grippers and the inclined perching object. For even larger inclinations, the rising axial momentum of the system becomes the driving factor and diminishes these gains again. Overall, the highest impact velocity leading to 100 % successful perching attempts, up to and including 9° of inclination, is 1.1 m/s.